### v2

@@ Abstract
A key challenge in scientific machine learning is solving partial differential equations (PDEs) on complex domains, where the curved geometry complicates the approximation of functions and their derivatives required by differential operators. This paper establishes the first simultaneous approximation theory for deep neural networks on $d$-dimensional manifolds $\mathcal{M}^{d}$. We prove that a constant-depth $\mathrm{ReLU}^{k-1}$ network with bounded weights—a property that plays a crucial role in controlling generalization error—can approximate any function in the Sobolev space $\mathcal{W}_{p}^{k}(\mathcal{M}^{d})$ to an error of $\varepsilon$ in the $\mathcal{W}_{p}^{s}(\mathcal{M}^{d})$ norm, for $k\geq 3$ and $s<k$, using $\mathcal{O}(\varepsilon^{-d/(k-s)})$ nonzero parameters, a rate that overcomes the curse of dimensionality by depending only on the intrinsic dimension $d$. These results readily extend to functions in Hölder-Zygmund spaces. We complement this result with a matching lower bound, proving our construction is nearly optimal by showing that the required number of parameters matches up to a logarithmic factor. Our proof of the lower bound introduces novel estimates for the Vapnik-Chervonenkis dimension and pseudo-dimension of the network’s high-order derivative classes. These complexity bounds provide a theoretical cornerstone for learning PDEs on manifolds involving derivatives. Our analysis reveals that the network architecture leverages a sparse structure to efficiently exploit the manifold’s low-dimensional geometry.
@@ 1 Introduction
The remarkable success of deep learning in industrial applications, such as computer vision [32], natural language processing [20], and robotics [22], has spurred its growing adoption in scientific computing. While classical numerical methods—like finite element and spectral methods—possess well-established theoretical foundations for low-dimensional Partial Differential Equations (PDEs), they face two major obstacles: the curse of dimensionality in high-dimensional settings, and the significant computational challenges posed by complex geometries. To overcome these critical challenges, deep neural networks (DNNs) have emerged as powerful surrogate models by leveraging their strong nonlinear approximation capabilities. This has engendered novel numerical paradigms for tackling PDE problems, which can be broadly classified into two categories: unsupervised, physics-driven algorithms, such as Physics-Informed Neural Networks (PINNs) [39, 30, 10, 57], and supervised, data-driven algorithms like neural operators [36, 31]. For a comprehensive review of deep learning approaches for solving PDEs, we refer the reader to [15]. These advances highlight deep learning’s advantage in high-dimensional, nonlinear problems and open new avenues for modeling complex physical phenomena beyond the reach of traditional methods.
This empirical success is theoretically grounded in the universal approximation theorems of the 1980s and 1990s [13, 38, 26]. These seminal works established that even a simple, shallow network can approximate any continuous function to arbitrary precision. However, these classical theorems are purely qualitative. They guarantee the existence of an adequate network approximator (i.e., establishing class density) but are non-constructive: they offer no quantitative guidance on how network architecture, such as its size and depth, relates to a prescribed level of accuracy $\varepsilon$.
@@
Concurrently, theoretical work has begun to establish approximation guarantees for Hölder functions on manifolds. Leveraging tools like coordinate charts or random projections (i.e., the Johnson-Lindenstrauss lemma [17]), these studies achieve approximation rates dependent solely on the manifold’s intrinsic dimension $d$ [40, 12, 33]. Despite this progress, these results are insufficient for solving PDEs due to two key limitations. First, the error metrics are restricted to $L_{p}$ norms, which provide no control over derivatives—a prerequisite for handling operators such as the Laplace-Beltrami operator ($\Delta_{\mathcal{M}}$). Second, the Hölder spaces in use often depend on an ambient Euclidean metric [40, 33], making their definition non-intrinsic and complicating practical verification.
@@
In this work, we establish the approximation rates achieved by deep neural networks with respect to high-order Sobolev norms for functions defined on a broad class of $d$-dimensional Riemannian manifolds $\mathcal{M}^{d}$ with bounded geometry. Our approach marks a significant departure from prior work, which has largely focused on highly specialized settings. A notable example is a pioneering study on simultaneous approximation for the sphere [35]. This work established rates for convolutional neural networks (CNNs) by ingeniously exploiting the sphere’s unique zonal structure. On the sphere, the reproducing kernels of diffusion polynomial spaces depend only on the geodesic distance, which in turn is a simple function of the ambient Euclidean inner product. This specific geometric property was key to leveraging the known efficiency of CNNs in modeling functions of inner products [56, 55], thereby achieving simultaneous approximation. Our theory, in contrast, relies on no such specialized structural assumptions.
@@
The CNN-based methodology is not restricted to the sphere. Its underlying mechanism—the zonal structure—is a fundamental characteristic of two-point homogeneous spaces [48], a class distinguished by the existence of a convenient addition formula [19]. This class of manifolds, however, is exceptionally restrictive. It has been fully classified and consists of just a few canonical families: spheres, projective manifolds, and the Cayley projective plane. Our work takes a more general path, free from such constraints, with sparse localization constructions that may in turn illuminate the feasibility of CNN-like architectures beyond these special cases.
@@
To establish the optimality of these rates, we derive a matching lower bound for the Sobolev space on manifolds $\mathcal{W}^{k}_{\infty}(\mathcal{M}^{d})$. Specifically, we demonstrate that approximating any function in this space to an accuracy $\varepsilon$ (in the $\mathcal{W}^{s}_{\infty}(\mathcal{M}^{d})$ norm) necessitates a constant-depth $\mathrm{ReLU}^{k-1}$ network with at least $S$ parameters, where $S\,\log S\;\gtrsim\;\varepsilon^{-\frac{d}{\,k-s\,}}.$ This result confirms that the network complexity is essentially governed by the manifold’s intrinsic dimension $d$, not the ambient one. Our proof introduces novel estimates for both the Vapnik-Chervonenkis (VC) dimension and the pseudo-dimension of the function class realized by the $s$-th order derivatives of $\mathrm{ReLU}^{k-1}$ networks. These complexity bounds represent a contribution of independent interest. Whereas existing complexity analyses for $\mathrm{ReLU}$ networks have focused on the network’s output or its first derivatives [4, 51], our work provides the first such bounds for high-order derivative classes. This is a critical prerequisite for developing a complete convergence theory for PINNs governed by high-order PDEs—a gap our results directly fill.
@@
This paper is organized as follows. Section 2 introduces our assumptions, several equivalent definitions of Sobolev norms on manifolds, and the neural network architecture. Section 3 presents our main theorems, establishing matching upper and lower bounds for approximation rates in Sobolev and Hölder-Zygmund spaces, which demonstrates the optimality of the network complexity. We discuss the implications and future research directions in Section 4. Proofs of our main results are provided in Section 5, while all notation, auxiliary lemmas, and their proofs are collected in the Appendix.
@@ 2.1 Sobolev and Hölder-Zygmund Functions on Manifolds
Here, $CH^{s}(\mathbb{R}^{d})$ denotes the classical Hölder-Zygmund space on $\mathbb{R}^{d}$, which is equivalent to the Besov space $B^{s}_{\infty,\infty}$ (see [46, Section 1.2.2 and 1.5.1]). We consider the case where $s$ is not an integer, for which the Hölder-Zygmund space coincides with the classical Hölder space. Specifically, its norm is defined as:
@@
A proof of this equivalence can be found in [46, Sec. 1.2.2].
@@ 3 Main Results
Notably, the network complexity depends only on the manifold’s intrinsic dimension $d$. This demonstrates the ability of neural networks to overcome the curse of dimensionality, as their complexity scales with $d$ rather than the potentially much larger ambient dimension $D$. An additional result, presented as a corollary, extends this principle to the Hölder-Zygmund family of functions, showing that the network’s ability to exploit low-dimensional geometric structure holds for this function class as well.
@@
The proof strategy is inspired by [4, Theorem 6] and [28, Theorem 4.13]. The central step involves reconstructing each function in the derivative class $D^{\alpha}\mathcal{F}$ with a new, surrogate network. This network employs a combination of activation functions, namely $\mathrm{ReLU}^{t}$ for $t=s,s+1,\ldots,k-1$. This key constructive argument is rigorously presented in Lemma 5. Consequently, the problem reduces to estimating the complexity of this new function class.

### v3

@@ Abstract
A key challenge in scientific machine learning is solving partial differential equations (PDEs) on complex domains, where the curved geometry complicates the approximation of functions and their derivatives required by differential operators. This paper establishes the first simultaneous approximation theory for deep neural networks on $d$-dimensional manifolds $\mathcal{M}^{d}$. We prove that a constant-depth $\mathrm{ReLU}^{k-1}$ network with bounded weights—a property that plays a crucial role in controlling generalization error—can approximate any function in the Sobolev space $\mathcal{W}_{p}^{k}(\mathcal{M}^{d})$ to an error of $\varepsilon$ in the $\mathcal{W}_{p}^{s}(\mathcal{M}^{d})$ norm, for $k\geq 3$ and $s<k$, using $\mathcal{O}(\varepsilon^{-d/(k-s)})$ nonzero parameters. This rate overcomes the curse of dimensionality by depending only on the intrinsic dimension $d$. These results readily extend to functions in Hölder–Zygmund spaces. We complement this result with a matching lower bound, proving our construction is nearly optimal by showing that the required number of parameters matches up to a logarithmic factor. Our proof of the lower bound introduces novel estimates for the Vapnik–Chervonenkis dimension and pseudo-dimension of the network’s high-order derivative classes. These complexity bounds provide a theoretical cornerstone for learning PDEs on manifolds involving derivatives. Our analysis reveals that the network architecture leverages a sparse structure to efficiently exploit the manifold’s low-dimensional geometry. Finally, we corroborate our theoretical findings with numerical experiments.
@@ 1 Introduction
The remarkable success of deep learning in industrial applications, such as computer vision [35], natural language processing [21], and robotics [23], has spurred its growing adoption in scientific computing. While classical numerical methods—like finite element and spectral methods—possess well-established theoretical foundations for low-dimensional Partial Differential Equations (PDEs), they face two major obstacles: the curse of dimensionality in high-dimensional settings, and the significant computational challenges posed by complex geometries. To overcome these critical challenges, deep neural networks (DNNs) have emerged as powerful surrogate models by leveraging their strong nonlinear approximation capabilities. This has engendered novel numerical paradigms for tackling PDE problems, which can be broadly classified into two categories: unsupervised, physics-driven algorithms, such as physics-informed neural networks (PINNs) [42, 32, 10, 59], and supervised, data-driven algorithms like neural operators [39, 34]. For a comprehensive review of deep learning approaches for solving PDEs, we refer the reader to [16]. These advances highlight the advantage of deep learning in high-dimensional, nonlinear problems and open new avenues for modeling complex physical phenomena beyond the reach of traditional methods.
This empirical success is theoretically grounded in the universal approximation theorems of the 1980s and 1990s [13, 41, 28]. These seminal works established that even a simple, shallow network can approximate any continuous function to arbitrary precision. However, these classical theorems are purely qualitative. They guarantee the existence of a suitable network approximator (i.e., they establish class density), but the results are non-constructive: they provide no quantitative guidance on how network architecture—such as its width or depth—should scale with a prescribed accuracy level $\varepsilon$.
@@
Concurrently, theoretical work has begun to establish approximation guarantees for Hölder functions on manifolds. Leveraging tools like coordinate charts or random projections (i.e., the Johnson-Lindenstrauss lemma [18]), these studies achieve approximation rates dependent solely on the manifold’s intrinsic dimension $d$ [43, 12, 36]. Despite this progress, these results are insufficient for solving PDEs due to two key limitations. First, the error metrics are restricted to $L_{p}$ norms, which provide no control over derivatives—a prerequisite for handling operators such as the Laplace-Beltrami operator $\Delta_{\mathcal{M}}$. Second, the Hölder spaces in use often depend on an ambient Euclidean metric [43, 36], making their definition non-intrinsic and complicating practical verification.
@@
In this work, we establish the approximation rates achieved by deep neural networks with respect to high-order Sobolev norms for functions defined on a broad class of $d$-dimensional Riemannian manifolds $\mathcal{M}^{d}$ with bounded geometry. Our approach marks a significant departure from prior work, which has primarily focused on highly specialized settings. A notable example is a pioneering study on simultaneous approximation for the sphere [38]. This work established rates for convolutional neural networks (CNNs) by ingeniously exploiting the sphere’s unique zonal structure. On the sphere, the reproducing kernels of diffusion polynomial spaces depend only on the geodesic distance, which in turn is a simple function of the ambient Euclidean inner product. This specific geometric property was key to leveraging the known efficiency of CNNs in modeling functions of inner products [58, 57], thereby achieving simultaneous approximation. Our theory, in contrast, relies on no such specialized structural assumptions.
@@
The CNN-based methodology is not limited to a specific sphere. Its underlying mechanism—the zonal structure—is a fundamental characteristic of two-point homogeneous spaces [51], a class distinguished by the existence of a convenient addition formula [20]. This class of manifolds, however, is exceptionally restrictive. It has been fully classified and consists of just a few canonical families: spheres, projective manifolds, and the Cayley projective plane. Our work takes a more general approach, free from such constraints, with sparse localization constructions that may in turn illuminate the feasibility of CNN-like architectures beyond these special cases.
@@
To establish the optimality of these rates, we derive a matching lower bound for the Sobolev space on manifolds $\mathcal{W}^{k}_{\infty}(\mathcal{M}^{d})$. Specifically, we demonstrate that approximating any function in this space to an accuracy $\varepsilon$ (in the $\mathcal{W}^{s}_{\infty}(\mathcal{M}^{d})$ norm) necessitates a constant-depth $\mathrm{ReLU}^{k-1}$ network with at least $S$ parameters, where $S\,\log S\;\gtrsim\;\varepsilon^{-\frac{d}{\,k-s\,}}.$ This result confirms that the network complexity is essentially governed by the manifold’s intrinsic dimension $d$, not the ambient one. Our proof introduces novel estimates for both the Vapnik-Chervonenkis (VC) dimension and the pseudo-dimension of the function class realized by the $s$-th order derivatives of $\mathrm{ReLU}^{k-1}$ networks. These complexity bounds represent a contribution of independent interest. Whereas existing complexity analyses for $\mathrm{ReLU}$ networks have focused on the network’s output or its first derivatives [4, 54], our work provides the first such bounds for high-order derivative classes. This is a critical prerequisite for developing a complete convergence theory for PINNs governed by high-order PDEs, which our results address.
@@
This paper is organized as follows. Section 2 introduces our assumptions, several equivalent definitions of Sobolev norms on manifolds, and the neural network architecture. Section 3 presents our main theorems, establishing matching upper and lower bounds for approximation rates in Sobolev and Hölder–Zygmund spaces, which demonstrates the optimality of the network complexity. In Section 4 we give numerical experiments that validate the theoretical findings and provide architectural design insights. We discuss the implications and future research directions in Section 5. Proofs of our main results are provided in Section 6, while all notation, auxiliary lemmas, and their proofs are collected in the Appendix.
@@ 2.1 Sobolev and Hölder-Zygmund Functions on Manifolds
Here, $CH^{s}(\mathbb{R}^{d})$ denotes the classical Hölder-Zygmund space on $\mathbb{R}^{d}$, which is equivalent to the Besov space $B^{s}_{\infty,\infty}$ (see [49, Section 1.2.2 and 1.5.1]). We consider the case where $s$ is not an integer, for which the Hölder-Zygmund space coincides with the classical Hölder space. Specifically, its norm is defined as:
@@
A proof of this equivalence can be found in [49, Sec. 1.2.2].
@@ 3 Main Results
Notably, the network complexity depends only on the manifold’s intrinsic dimension $d$. This demonstrates the ability of neural networks to overcome the curse of dimensionality, as their complexity scales with $d$ rather than the potentially much larger ambient dimension $D$. An additional result, presented as a corollary, extends this principle to the Hölder-Zygmund family of functions, demonstrating that the network’s ability to exploit low-dimensional geometric structure also applies to this function class.
@@
The proof strategy is inspired by [4, Theorem 6] and [30, Theorem 4.13]. The central step involves reconstructing each function in the derivative class $D^{\alpha}\mathcal{F}$ with a new, surrogate network. This network employs a combination of activation functions, namely $\mathrm{ReLU}^{t}$ for $t=s,s+1,\ldots,k-1$. This key constructive argument is rigorously presented in Lemma 5. Consequently, the problem reduces to estimating the complexity of this new function class.
@@
4 Experiments
In this section, we conduct numerical experiments to validate our theoretical findings and provide architectural design insights. The experiments are implemented using Python scripts powered by the PyTorch framework https://pytorch.org/. The scripts can be downloaded from https://github.com/hanfei27/ReLUkNNonManifolds.
4.1 Numerical Experiment Setup
Denote the unit sphere by $\mathbb{S}^{2}=\{x\in\mathbb{R}^{3}:\|x\|=1\}.$ On $\mathbb{S}^{2}$ the target function is the standardly normalized third-order zonal spherical harmonic
$f(x)=Y_{3,1}(x),\qquad x\in\mathbb{S}^{2},$
which satisfies the spectral relation $\Delta_{\mathbb{S}^{2}}Y_{3,1}=-\lambda_{3}Y_{3,1}$ with $\lambda_{3}=12$. Definitions and the relevant spectral properties of the Laplace–Beltrami operator $\Delta_{\mathbb{S}^{2}}$ and of spherical harmonics are collected in Appendix C.
We generate a weighted sample set $\{(x_{i},w_{i})\}_{i=1}^{N}$ on $\mathbb{S}^{2}$ using a Fibonacci grid [26] and approximate surface integrals by the discrete quadrature
$\int_{\mathbb{S}^{2}}g(x)\,\mathrm{d}\omega(x)\approx\sum_{i=1}^{N}w_{i}\,g(x_{i}).$
The model is a fully connected $\mathrm{ReLU}^{k}$ network $u_{\theta}:\mathbb{R}^{3}\to\mathbb{R}$, evaluated at the sampled points. During training we optimise only the $L^{2}$ (function value) discrepancy; explicitly the training objective on the sphere is
$\mathcal{L}_{\mathbb{S}^{2}}^{\mathrm{train}}(\theta)=\sum_{i=1}^{N}w_{i}\big|u_{\theta}(x_{i})-f(x_{i})\big|^{2}.$
Training uses the Adam optimizer [33] with initial learning rate $\eta=10^{-3}$, batch size $\mathrm{BS}=2048$, and runs for $T$ iterations. Optimization is performed by stochastic minibatch updates over the weighted samples and the best parameters found during training are saved for evaluation.
For evaluation we measure Sobolev-style errors that probe first- and second-order accuracy, even though these terms are not included in the training objective. Concretely we compute the weighted mean squared errors
$\mathrm{WMSE}_{f,\mathbb{S}^{2}}$ $=\sum_{i=1}^{N}w_{i}\big|u_{\theta}(x_{i})-f(x_{i})\big|^{2},$
$\mathrm{WMSE}_{\nabla,\mathbb{S}^{2}}$ $=\sum_{i=1}^{N}w_{i}\big\|\nabla_{\mathbb{S}^{2}}u_{\theta}(x_{i})-\nabla_{\mathbb{S}^{2}}f(x_{i})\big\|^{2},$
$\mathrm{WMSE}_{\Delta,\mathbb{S}^{2}}$ $=\sum_{i=1}^{N}w_{i}\big|\Delta_{\mathbb{S}^{2}}u_{\theta}(x_{i})-\Delta_{\mathbb{S}^{2}}f(x_{i})\big|^{2}.$
The tangential gradients and the Laplace–Beltrami operator used in these test-time metrics are computed numerically as described in Appendix C. Spectrally, the gradient and Laplace metrics correspond to multiplying spherical-harmonic coefficients by $\lambda_{n}$ and $\lambda_{n}^{2}$, respectively, so these weighted MSEs probe approximation quality across frequency bands and provide a Sobolev-style assessment of the learned model.
We repeat the validation on an embedded torus $\mathbb{T}^{2}_{R,r}\subset\mathbb{R}^{3}$ parameterized by
$X_{R,r}(u,v)=\big((R+r\cos v)\cos u,\;(R+r\cos v)\sin u,\;r\sin v\big),\qquad u,v\in[0,2\pi),$
with $R>r>0$ (in experiments we commonly take R=1.5,r=0.5R=1.5,\;r=0.5). The torus target is taken from an analytic Fourier family, for example
$g(x)=A\cos(m\,u(x))+B\sin(n\,v(x)),$
where $u(x),v(x)$ are the angular coordinates recovered from $x$. The Laplace–Beltrami action on this family admits a closed-form expression (see Appendix C), hence $\Delta_{\mathbb{T}^{2}}g$ is available exactly for testing.
Weighted torus samples $\{(y_{j},\tilde{w}_{j})\}_{j=1}^{M}$ are drawn by sampling $(u_{j},v_{j})$ uniformly on $0,2\pi)^{2}$, mapping via $X_{R,r}$, and rescaling weights by the Jacobian $\tilde{w}_{j}\propto r(R+r\cos v_{j})$ so that
$\int_{\mathbb{T}^{2}_{R,r}}h\,\mathrm{d}\tilde{\omega}(y)\approx\sum_{j=1}^{M}\tilde{w}_{j}\,h(y_{j}).$
On the torus we train with the $L^{2}$ objective
$\mathcal{L}_{\mathbb{T}^{2}}^{\mathrm{train}}(\theta)=\sum_{j=1}^{M}\tilde{w}_{j}\big|u_{\theta}(y_{j})-g(y_{j})\big|^{2},$
using the same Adam schedule as on the sphere. Performance is reported via $\mathrm{WMSE}_{g,\mathbb{T}^{2}}$, $\mathrm{WMSE}_{\nabla,\mathbb{T}^{2}}$ and $\mathrm{WMSE}_{\Delta,\mathbb{T}^{2}}$; under the toroidal Fourier basis these correspond spectrally to multiplying coefficients by $1$, $\mu_{m,n}$ and $\mu_{m,n}^{2}$, respectively, and thus quantify accuracy across the doubly periodic frequency bands consistent with the Sobolev norm on $\mathbb{T}^{2}_{R,r}$.
4.2 Network Architectural Ablation Study
We evaluated the effect of activation order by sweeping $\mathrm{act\_k\_list}=\{1,2,3,4,5,6,7\}$ (elementwise activation $\sigma_{k}(z)=(\mathrm{ReLU}(z))^{k}$; $k=1$ is standard ReLU). Experiments were performed on both the unit sphere $\mathbb{S}^{2}$ and the embedded torus $\mathbb{T}^{2}_{R,r}$; all runs used the identical baseline configuration given in the main text. Each $k$ was repeated 5 times on each manifold and Table 1 below reports the final training loss (mean $\pm$ std, 5 repeats) for both manifolds.
Table 1: Effect of activation order $k$ on final training loss (mean $\pm$ std, 5 repeats).
$k$ training loss (sphere) training loss (torus)
1 $\mathtt{9.42e-05}\pm\mathtt{2.70e-05}$ $\mathtt{3.22e-05}\pm\mathtt{4.85e-06}$
2 $\mathtt{4.19e-05}\pm\mathtt{3.99e-05}$ $\mathtt{1.48e-05}\pm\mathtt{6.96e-06}$
3 $\mathtt{5.23e-05}\pm\mathtt{1.85e-05}$ $\mathtt{1.64e-05}\pm\mathtt{1.52e-05}$
4 $\mathtt{9.75e-06}\pm\mathtt{2.11e-06}$ $\mathtt{6.75e-06}\pm\mathtt{4.78e-06}$
5 $\mathtt{8.45e-06}\pm\mathtt{2.65e-06}$ $\mathtt{3.35e-06}\pm\mathtt{2.02e-06}$
6 $\mathtt{4.36e-06}\pm\mathtt{1.42e-06}$ $\mathtt{4.40e-06}\pm\mathtt{3.75e-06}$
7 $\mathtt{2.82e-06}\pm\mathtt{7.09e-07}$ $\mathtt{3.68e-06}\pm\mathtt{1.22e-06}$
Additionally, we plot the function value, gradient, and Laplace–Beltrami loss components as functions of the activation order $k$. Solid lines indicate the sample mean across five repeats and shaded bands denote $\pm 1\sigma$. Results for both manifolds are shown: the unit sphere $\mathbb{S}^{2}$ and the embedded torus $\mathbb{T}^{2}_{R,r}$.
Figure 1: Component-wise mean MSEs (log10 scale) versus ReLU activation order $k$ on two manifolds. Left: unit sphere $\mathbb{S}^{2}$. Right: embedded torus $\mathbb{T}^{2}_{R,r}$. Plotted are function value MSE (blue), tangential gradient MSE (orange) and Laplace–Beltrami MSE (green). Curves denote the sample mean over five independent runs; shaded bands indicate $\pm 1\sigma$ and markers/error bars show the sample standard deviation.
In both the sphere and torus experiments we observe that the component MSEs (function value, tangential gradient and Laplace–Beltrami) decrease as the activation order $k$ increases (plots are shown on a log scale). This behavior aligns with two complementary theoretical points. First, standard $\mathrm{ReLU}$ networks (kk=1) are consistent with the simultaneous-approximation results of [24]: they can achieve control of the function and its first derivatives, i.e. first-order simultaneous approximation. Second, as we show in our analysis, ReLU networks do not, in general, provide simultaneous approximation beyond first order and therefore cannot guarantee uniform control of second-order quantities such as the Laplace–Beltrami. Increasing the activation order mitigates this limitation because higher-order activations are smoother at the kink and thus enable the network to form higher-degree piecewise-polynomial, spline-like approximants. This additional smoothness enables network exploit higher-order information of the target and assemble higher-degree local polynomial pieces, which improves approximation in Sobolev norms that are sensitive to second derivatives; accordingly, the Laplace–Beltrami MSE decreases as $k$ increases in our experiments. To illustrate this effect, Figure 2 shows pointwise spatial error fields (function value error, tangential gradient norm error, and absolute Laplace error) for $k=1$ and $k=5$ on both manifolds, highlighting the smoothing effect and local reductions in gradient / Laplace errors produced by the higher-order activation.
Figure 2: Spatial error maps on the sphere and torus.
Having characterized the effect of activation order $k$ and observed diminishing returns for large $k$, we henceforth fix $k=4$ and examine network width $m\in\{16,32,64,96,128,256\}$ (depth $L=2$), holding all other hyperparameters constant. The network depth was fixed at 2 so that width is the only variable. All runs used the same baseline configuration: total training steps $T=5000$, total samples $N=20000$, mini-batch size $\mathrm{BS}=2048$, learning rate $\eta=10^{-3}$, and random seed fixed for repeatability. Each width was trained independently for 5 repeats; reported statistics are the sample mean and sample standard deviation across repeats. The table below lists the three component MSEs (function value, gradient, Laplace–Beltrami).
Table 2: Final training loss (mean $\pm$ std, 5 repeats) for different network widths.
width $m$ training loss (sphere) training loss (torus)
16 $\mathtt{1.07e-03}\pm\mathtt{1.66e-03}$ $\mathtt{1.01e-03}\pm\mathtt{2.14e-04}$
32 $\mathtt{5.69e-05}\pm\mathtt{2.47e-05}$ $\mathtt{2.42e-04}\pm\mathtt{5.26e-05}$
64 $\mathtt{2.12e-05}\pm\mathtt{1.54e-05}$ $\mathtt{8.92e-05}\pm\mathtt{2.39e-05}$
96 $\mathtt{1.80e-05}\pm\mathtt{1.26e-05}$ $\mathtt{7.09e-05}\pm\mathtt{1.63e-05}$
128 $\mathtt{5.18e-06}\pm\mathtt{2.42e-06}$ $\mathtt{5.46e-05}\pm\mathtt{1.22e-05}$
256 $\mathtt{1.81e-06}\pm\mathtt{3.80e-07}$ $\mathtt{4.47e-05}\pm\mathtt{1.07e-05}$
We also plot the component-wise MSEs in two log–log diagrams that both display MSE versus network width: one subplot for the sphere and one for the torus (horizontal axis: $\log_{2}(\text{width})$). Each curve shows the sample mean across five repeats with shaded bands indicating $\pm 1\sigma$.
Figure 3: Component-wise mean MSEs versus network width (log–log). Left: unit sphere $\mathbb{S}^{2}$. Right: embedded torus $\mathbb{T}^{2}_{R,r}$. Plotted are function value MSE (blue), tangential gradient MSE (orange) and Laplace–Beltrami MSE (green). Solid lines denote the sample mean over five independent runs; shaded bands indicate $\pm 1\sigma$. The horizontal axis is shown as $\log_{2}(\text{width})$ and the vertical axis as $\log_{10}(\text{MSE})$.
Both plots (sphere and torus) display an approximately linear decline on log–log axes, so we fit simple linear models in the logarithmic domain to estimate power-law decay exponents for each error component. The fitted lines indicate algebraic (power-law) decay of the error with model size (width / parameter count) in both cases, which qualitatively supports the mechanism predicted by our theory. Quantitatively, however, the fitted slopes do not match the theoretical worst-case exponents: the theory provides conservative bounds valid in the worst case, whereas our numerical targets are much smoother than worst-case examples — the sphere experiments use a spherical harmonic and the torus experiments use Fourier-type targets, both of which possess strong regularity. Consequently exact agreement of exponents is not expected, although the observed algebraic decay corroborates the theory’s qualitative prediction that larger models yield polynomially improving Sobolev accuracy. The fitted log–log slopes for each error component (fit of $\log_{10}(\mathrm{MSE})$ versus $\log_{10}(\text{parameter count})$) are reported in  Table 3. The slopes quantify the observed algebraic decay rates; the sphere exhibits substantially steeper decay than the torus, reflecting the higher regularity of the spherical-harmonic target used in those runs.
Table 3: Estimated log–log slopes $\alpha$ (fit of $\log_{10}(\mathrm{MSE})$ vs. $\log_{10}$ parameter count).
Component Sphere (log–log slope) Torus (log–log slope)
Value 2.1708-2.1708 1.1323-1.1323
Grad 1.9832-1.9832 1.1741-1.1741
Laplace 1.7896-1.7896 1.2026-1.2026
We also studied network depth. Empirically, for activation order $k>3$ increasing depth tends to destabilize optimization and can prevent convergence. To mitigate this for high-order $\mathrm{ReLU}^{k}$ in deep nets we applied a systematic stabilization protocol: (i) a variance rescaling based on half-normal moments so that
$\operatorname{Var}\big(s_{k}\mathrm{ReLU}^{\,k}(z)\big)=\operatorname{Var}\big(\mathrm{ReLU}(z)\big),$
with scalar $s_{k}>0$ estimated from the half-normal assumption on pre-activations; (ii) insertion of LayerNorm after each hidden linear layer; (iii) upper bounds on activations and weights ($\text{activation\_clamp\_max}=3.0,\ \text{weight\_clip}=0.1$); and (iv) an optimization schedule with reduced initial learning rate $\eta=5\times 10^{-4}$. With these measures a $\mathrm{ReLU}^{k}$ network with $k=5$ and depth $L=5$ stably converges on the spherical-harmonic target, reducing the Laplace–Beltrami residual to $\mathcal{O}(10^{-2})$. However, the deep model shows substantially longer training time and no clear error advantage over the shallower $L=2$ baseline; this empirical finding aligns with our theoretical message that, absent explicit parameter-norm constraints, modest depth (e.g. $L=3$) already suffices to obtain the predicted approximation rates.
4.3 Mixed Architectures Study
We assessed whether $\mathrm{ReLU}$$\mathrm{ReLU}^{k}$ mixtures can combine the optimization advantages of ReLU with the higher-order approximation power of smooth activations. Experimental settings were held fixed: network width $m=128$, depth $L=4$, training samples N=10, [^0] N=10{,}000, training steps T=4, [^0] T=4{,}000, and LayerNorm applied uniformly. The activation pattern was the only variable: a tail-mixed family $[1,1,1,5]$ (first three hidden layers $\mathrm{ReLU}$, final hidden layer $\mathrm{ReLU}^{5}$) and single high-order placements $[5,1,1,1]$, $[1,5,1,1]$, $[1,1,5,1]$ were tested. To exclude simple optimizer remedies, we ran ablations with activation clamping on and off, as well as with several gradient-clipping thresholds.
Across all configurations and optimization variants, the weighted Laplace error remained on the order of $\sim 4.4\times 10^{1}$ to $\sim 1.25\times 10^{2}$, comparable to the first-order error level of standard ReLU networks; the corresponding gradient error likewise stayed at $O(10^{-2})$ with no systematic improvement. These results indicate that, under the present experimental conditions, ReLU–ReLUk mixed architectures (regardless of high-order unit placement) do not induce the expected high-order ($H^{2}$-type) convergence. In principle, a $\mathrm{ReLU}$$\mathrm{ReLU}^{k}$ mixed architecture can realize the smooth approximants required for $H^{2}$-type convergence; however, in practice the optimization process offers no guarantee that the $\mathrm{ReLU}$ components will be driven to the specific smooth configurations needed, so the theoretical feasibility does not translate into a reliable practical route to second-order convergence. Consequently, for improved numerical performance when second-order control is important, we recommend using globally smoother activations (for example, GeLU or softplus) or, as a pragmatic compromise that preserves gradient flow, LeakyReLU; these alternatives typically yield more stable optimization and better empirical reduction of Laplace-sensitive errors than architectures that mix non-smooth ReLU blocks with high-order units. However, a full exploration of these alternative activations and their theoretical properties lies beyond the scope of this paper.